OpenStudy (thefluffmuncher):

Which is a counterexample that disproves the conjecture? If n is a positive integer, then 2n – 1 is a prime number. A. n = 6 B. n = 5 C. n = 3 D. n = 2

OpenStudy (thefluffmuncher):

@rishi_84

ganeshie8 (ganeshie8):

prime numbers wud be very boring if that conjecture was true lol

ganeshie8 (ganeshie8):

plugin each option into the given expression and see which one does not give u a prime number...

OpenStudy (thefluffmuncher):

a is eliminated

OpenStudy (thefluffmuncher):

it's b right

ganeshie8 (ganeshie8):

lets test first option : A. n = 6 2n - 1 = 2(6) - 1 = 12 - 1 = 11 is a prime number, so its not a counter example. Yes eliminate this.

ganeshie8 (ganeshie8):

You got it ! B is a counter example cuz 2(5) - 1 = 10 - 1 = 9 is NOT a prime number

OpenStudy (thefluffmuncher):

yay im like half asleep and i got it right woohoo!

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

bro b will be option because 2*5-1 = 9 is not a prime